# Technical report on label-informed logit redistribution for better domain generalization in low-shot classification with foundation models

### v2

@@ A.1 Relation to Perplexity
A.2 Relation to Confidence Calibration in Deep Models
@@
$s(x,y^{\prime})\gg s(x,y_{j}),\quad\forall j\neq y^{\prime}.$
@@
$P(y^{\prime}|x;\theta)\approx 1,\quad P(y_{j}|x;\theta)\approx 0,\quad\forall j\neq y^{\prime}.$
@@
As $P(y^{\prime}|x;\theta)\to 1$, the entropy $H(P(y|x;\theta))\to 0$, confirming that the softmax output becomes overly confident.
@@
Appendix B Experiments and Results
@@
B.1 Implementation detail
Prompt learning baselines. In prompting baseline of our paper we run the baseline publicly available at their GitHub. However we modified the code for integrating penalty and then run the experiment.
Benchmarks. To check the effectiveness of RccPL in domain generalization settings, we implement RccPL integration into CLIP with ResNet-50 backbone while the results for CLIPCalib are reproduced using their publicly available code. We then performed evaluation on domain generalization datasets with $\lambda$ = $0.01$.
@@ B.1 Implementation detail
B.2 Results
RccPL performance in comparison with CLIPCalib with (ResNet-50) are presented in Table 5. It has been observed that CLIP baseline achieve moderate calibration with an average ECE of 7.8 in domain generalizations datasets. However it performs well in dataset Real (ECE 0.98) but poorly calibrated results on Sketch and Art (3.17 and 2.13).
@@ B.2 Results
Our method CLIPRccPL shows better performance by achieving lowest average ECE 7.32. RccPL surpasses on both baseline and benchmark on the Real (R), Sketch (S), and Art (A) datasets (0.94, 6.92, and 18.9, respectively). These results indicates that RccPL effectively improves in calibration in challenging domain generalization setting.
Table 5: RccPL performance results in comparison with CLIPCalib with (ResNet-50) as backbone for domain genralisation .
@@
CLIPRccPL 3.64 0.94 6.92 18.9 2.53 7.32
Table 6: Calibration performance of RccPL in comparison with all other prompt learning methods across individual datasets for different methods. The last column shows the average across all datasets for each method. ($\downarrow$) Represents that smaller values are better. ($\uparrow$) represents the larger values are better. Bold values indicate the best results.
@@
KgCoOp + (RccPL) 76.3 76.0 76.1 76.4 76.2 76.2 76.4 76.3 76.2 76.1 76.3 76.3
Table 7: RccPL $\lambda$ strength.
@@
$\lambda$ value CLIP CoOP + (RccPL) CoCoOp + (RccPL) ProDA + (RccPL) MaPLe + (RccPL) KgCoOp + (RccPL)

### v3

@@ B.1 Relation to Perplexity
B.2 Relation to Confidence Calibration in Deep Models
@@
$s(x,y)\gg s(x,y^{\prime}),\quad\forall j\neq y^{\prime}.$
@@
$P(y|x;\theta)\approx 1,\quad P(y^{\prime}|x;\theta)\approx 0,\quad\forall j\neq y^{\prime}.$
@@
As $P(y|x;\theta)\to 1$, the entropy $H(P(y|x;\theta))\to 0$, confirming that the softmax output becomes overly confident.
@@
Appendix C Experiments and Results
@@
C.1 Implementation detail
Prompt learning baselines. In prompting baseline of our paper we run the baseline publicly available at their GitHub. We run the code as is, except for integrating penalty and then conduct the experiment.
C.2 Reproducibility
We release code, pre-trained prompts, and evaluation scripts at given anonymous link cmp. to facilitate full reproducibility of results.
Benchmarks. To check the effectiveness of CMP in domain generalization settings, we implement CMP integration into CLIP with ResNet-50 backbone while the results for CLIPCalib are reproduced using their publicly available code. We then performed evaluation on domain generalization datasets with $\lambda$ = $0.01$.
@@ C.2 Reproducibility
C.3 Results
CMP performance in comparison with CLIPCalib with (ResNet-50) are presented in Table 5. It has been observed that CLIP baseline achieve moderate calibration with an average ECE of 7.8 in domain generalizations datasets. However it performs well in dataset Real (ECE 0.98) but poorly calibrated results on Sketch and Art (3.17 and 2.13).
@@ C.3 Results
Our method CLIPCMP shows better performance by achieving lowest average ECE 7.32. CMP surpasses on both baseline and benchmark on the Real (R), Sketch (S), and Art (A) datasets (0.94, 6.92, and 18.9, respectively). These results indicates that CMP effectively improves in calibration in challenging domain generalization setting.
Table 5: CMP performance results in comparison with CLIPCalib with (ResNet-50) as backbone for domain generalisation.
@@
C-tpt 2.42 1.45 6.85 22.1 1.75 8.51
B-PEFT 2.40 1.63 6.98 22.9 1.78 8.74
CLIPCMP 3.64 0.94 6.92 18.9 2.53 7.32
Table 6: Calibration performance of CMP across few-shot settings (1/2/4-shot). Bold values represent significant performance. In the left column Acc $\uparrow$ demonstrates larger value is significant. In the right column ECE $\downarrow$ shows smaller value is significant.
Method Acc($\uparrow$) ACE(↓) MCE(↓) ECE(↓)
0-shot
All methods 72.7 3.51 0.89 3.47
1-Shot
CLIP 72.7 3.51 0.89 3.47
CoOp 78.5 3.45 1.21 3.32
CoOp + CMP 79.1 3.18 1.05 3.09
CoCoOp 74.2 3.72 0.98 3.65
CoCoOp + CMP 73.8 3.68 0.94 3.61
ProDA 78.9 4.35 1.38 4.21
ProDA + CMP 79.5 4.02 1.32 3.95
MaPLe 78.2 4.71 1.52 3.25
MaPLe + CMP 79.8 4.15 1.28 2.95
KgCoOp 75.6 4.69 1.14 4.31
KgCoOp + CMP 76.1 4.35 1.02 4.12
2-Shot
CLIP 73.1 3.48 0.87 3.42
CoOp 79.2 3.32 1.12 3.18
CoOp + CMP 80.0 3.05 0.98 2.95
CoCoOp 75.1 3.61 0.92 3.52
CoCoOp + CMP 75.8 3.48 0.88 3.38
ProDA 79.7 4.22 1.29 4.05
ProDA + CMP 80.3 3.88 1.24 3.78
MaPLe 79.0 4.58 1.43 3.12
MaPLe + CMP 80.7 3.98 1.21 2.81
KgCoOp 76.5 4.55 1.08 4.15
KgCoOp + CMP 77.3 4.22 0.97 3.95
4-Shot
CLIP 73.5 3.45 0.85 3.38
CoOp 80.1 3.21 1.03 3.05
CoOp + CMP 81.2 2.94 0.91 2.82
CoCoOp 76.3 3.52 0.87 3.41
CoCoOp + CMP 77.1 3.38 0.82 3.25
ProDA 80.5 4.11 1.21 3.92
ProDA + CMP 81.4 3.75 1.15 3.62
MaPLe 80.2 4.42 1.35 2.98
MaPLe + CMP 81.9 3.82 1.12 2.65
KgCoOp 77.8 4.42 1.02 4.02
KgCoOp + CMP 78.9 4.08 0.93 3.81
Table 7: Comparison of CMP when combined with prompt learning methods across few-shot settings. Bold values represent significant performance. In the left column Acc $\uparrow$ demonstrates larger value is significant. In the right column ECE $\downarrow$ shows smaller value is significant.
Acc ($\uparrow$) ECE ($\downarrow$)
Method source target source target
ImageNet R S A V2 Avg ImageNet R S A V2 Avg
0-Shot
All methods 65.8 70.5 47.1 48.8 61.5 57.2 1.94 4.12 4.98 7.92 3.01 5.01
1-Shot
CLIP 65.8 70.5 47.1 48.8 61.5 57.2 1.94 4.12 4.98 7.92 3.01 5.01
CoOp 67.5 71.3 46.5 48.5 62.9 57.3 1.72 7.82 14.6 7.12 3.18 8.08
CoOp + CMP 68.2 72.1 46.9 49.1 63.5 57.9 1.65 7.15 13.8 6.65 2.45 7.42
CoCoOp 68.3 72.9 48.2 50.3 63.2 58.7 1.38 7.22 13.1 5.68 2.98 7.27
CoCoOp + CMP 69.2 74.1 48.7 51.2 63.9 59.5 1.45 6.95 12.2 5.05 2.15 6.56
ProDA 66.8 69.5 51.9 52.4 64.2 59.5 1.65 7.85 14.8 7.25 3.28 8.33
ProDA + CMP 65.7 67.8 51.5 51.5 64.5 58.8 1.58 6.45 14.0 7.35 3.05 7.68
MaPLe 69.9 70.8 52.5 48.4 62.9 58.7 1.78 4.05 6.25 14.1 2.28 6.92
MaPLe + CMP 68.3 71.2 53.1 50.5 63.5 59.6 1.85 3.05 4.65 13.2 2.18 5.83
KgCoOp 71.2 70.8 48.2 48.8 62.1 57.5 2.78 4.42 7.92 14.6 4.02 7.95
KgCoOp + CMP 70.5 72.1 47.5 49.9 63.3 58.2 2.95 4.35 7.45 13.8 3.35 7.38
2-Shot
CLIP 65.8 70.5 47.1 48.8 61.5 57.2 1.94 4.12 4.98 7.92 3.01 5.01
CoOp 68.3 72.1 46.8 49.2 63.8 58.0 1.58 7.65 14.3 6.85 3.05 7.96
CoOp + CMP 69.1 72.9 47.2 49.8 64.5 58.6 1.61 7.02 13.5 6.42 2.25 7.30
CoCoOp 69.2 73.8 49.1 51.2 64.1 59.6 1.25 7.05 12.8 5.45 2.85 7.05
CoCoOp + CMP 70.1 75.3 49.8 52.1 64.8 60.5 1.42 6.88 11.9 4.85 1.95 6.27
ProDA 67.9 70.2 52.8 53.1 65.1 60.3 1.52 7.68 14.5 7.05 3.15 8.10
ProDA + CMP 66.5 68.5 52.3 52.2 65.3 59.6 1.55 6.25 13.8 7.12 2.85 7.51
MaPLe 70.8 71.5 53.2 49.1 63.8 59.4 1.65 3.85 5.95 13.8 2.15 6.44
MaPLe + CMP 69.1 72.0 53.8 51.2 64.3 60.3 1.72 2.75 4.35 12.8 2.05 5.48
KgCoOp 72.1 71.5 48.9 49.5 62.9 58.2 2.65 4.25 7.65 14.3 3.85 7.51
KgCoOp + CMP 71.3 72.8 48.2 50.8 64.1 59.0 2.85 4.18 7.25 13.6 3.15 7.05
4-Shot
CLIP 65.8 70.5 47.1 48.8 61.5 57.2 1.94 4.12 4.98 7.92 3.01 5.01
CoOp 69.8 73.5 47.5 50.1 65.2 59.1 1.45 7.25 13.8 6.35 2.85 7.34
CoOp + CMP 70.9 74.2 48.1 50.8 66.1 59.8 1.38 6.55 12.9 5.92 2.05 6.76
CoCoOp 70.5 75.1 50.2 52.5 65.8 61.0 1.15 6.75 12.1 4.95 2.55 6.48
CoCoOp + CMP 71.8 76.5 50.9 53.2 66.5 61.7 1.25 6.35 11.2 4.35 1.75 5.98
ProDA 69.2 71.5 53.5 54.2 66.8 61.5 1.38 7.25 14.1 6.55 2.95 7.65
ProDA + CMP 67.8 69.8 53.0 53.5 67.2 60.9 1.42 5.85 13.2 6.65 2.65 7.15
MaPLe 71.5 72.8 54.1 50.2 65.2 60.6 1.52 3.55 5.45 13.2 1.95 5.93
MaPLe + CMP 70.2 73.5 54.8 52.5 65.9 61.7 1.58 2.45 3.95 12.1 1.85 5.19
KgCoOp 73.5 72.8 49.8 50.5 64.2 59.3 2.45 4.05 7.25 13.8 3.55 7.02
KgCoOp + CMP 72.8 74.2 49.2 51.8 65.5 60.2 2.65 3.95 6.85 13.1 3.05 6.72
Table 8: Calibration performance of CMP in comparison with all other prompt learning methods across individual datasets for different methods. The last column shows the average across all datasets for each method. ($\downarrow$) Represents that smaller values are better. ($\uparrow$) represents the larger values are better. Bold values indicate the best results.
@@
KgCoOp + (CMP) 76.3 76.0 76.1 76.4 76.2 76.2 76.4 76.3 76.2 76.1 76.3 76.3
Table 9: CMP $\lambda$ strength.
@@
$\lambda$ value CLIP CoOP + (CMP) CoCoOp + (CMP) ProDA + (CMP) MaPLe + (CMP) KgCoOp + (CMP)
@@
C.4 CMP Standalone Contribution
To isolate CMP’s effect, we compare standard CLIP fine-tuning with and without CMP on ImageNet (8-shot, ViT-B/16). Table 10 shows CMP alone improves accuracy by 1.8% while reducing calibration errors by 19.7% (ECE) and 36.0% (MCE), demonstrating its intrinsic value beyond prompt learning integration.
Table 10: CMP ablation on ImageNet (ViT-B/16, 8-shot)
Method Accuracy $\uparrow$ ECE $\downarrow$ ACE $\downarrow$ MCE $\downarrow$
CLIP (FT only) 82.3 4.12 3.51 1.89
CLIP (FT + CMP) 84.1 3.31 3.01 1.21
All hyperparameters remained identical between both setups. The consistent improvements across metrics confirm CMP’s standalone efficacy.